a plane travels 1600 miles against the wind in 10 hours and makes the return trip with the same tailwind in 8 hours. Find the rate of the plane
Answer by josgarithmetic(39626)

speed with wind r+c
speed against wind r-c
r is plane speed when no wind.

Rate*Time=Distance

(r-c)10=1600 and (r+c)8=1600
-
r-c=160
and
r+c=200

Use Elimination Method.
(r-c)+(r+c)=360
2r=360
r=180
and you can find c...